# SPIRAL: Self-Play on Zero-Sum Games Incentivizes Reasoning via Multi-Agent Multi-Turn Reinforcement Learning

### v2

@@ Abstract
Recent advances in reinforcement learning have shown that language models can develop sophisticated reasoning through training on tasks with verifiable rewards, but these approaches depend on human-curated problem-answer pairs and domain-specific reward engineering. We introduce SPIRAL, a self-play framework where models learn by playing multi-turn, zero-sum games against continuously improving versions of themselves, eliminating the need for human supervision. Through self-play, SPIRAL generates an infinite curriculum of progressively challenging problems as models must constantly adapt to stronger opponents. To enable this self-play training at scale, We implement a fully online, multi-turn, multi-agent reinforcement learning system for LLMs and propose role-conditioned advantage estimation (RAE) to stabilize multi-agent training. Using SPIRAL, self-play on zero-sum games produces reasoning capabilities that transfer broadly. Training Qwen3-4B-Base on Kuhn Poker alone achieves 8.6% improvement on math and 8.4% on general reasoning, outperforming SFT on 25,000 expert game trajectories. Analysis reveals that this transfer occurs through three cognitive patterns: systematic decomposition, expected value calculation, and case-by-case analysis. Multi-game training (TicTacToe, Kuhn Poker, Simple Negotiation) further enhances performance as each game develops distinct reasoning strengths. Applying SPIRAL to a strong reasoning model (DeepSeek-R1-Distill-Qwen-7B) can still lead to 2.0% average improvement. These results demonstrate that zero-sum games naturally develop transferable reasoning capabilities, highlighting a promising direction for autonomous reasoning development.                          [Uncaptioned image] Spiral-RL/SPIRAL       [Uncaptioned image] Spiral-RL
@@ 1 Introduction
However, current approaches to reasoning enhancement face a fundamental scalability bottleneck: their dependence on carefully engineered reward functions, domain-specific datasets, and expert supervision (DeepSeek Team, 2024; Ouyang et al., 2022; Bai et al., 2022). Each new reasoning domain requires experts to craft evaluation metrics, curate training problems, and validate reasoning traces. This manual process becomes increasingly unsustainable as we pursue more general intelligence, limiting both the scale and diversity of reasoning challenges that models can learn from.
@@
Self-play offers a solution to this scalability crisis by eliminating the need for human supervision in training data creation (Silver et al., 2017; Tesauro, 1995). In self-play, models learn by competing against copies of themselves, where each match outcome provides automatic feedback. As the model improves, its opponent improves equally, maintaining a consistent challenge level that drives continuous learning. This paradigm has already revolutionized AI across many domains: from TD-Gammon’s backgammon supremacy (Tesauro, 1995) to AlphaGo’s conquest of Go (Silver et al., 2016, 2017) to OpenAI Five’s mastery of complex team coordination (Berner et al., 2019). Yet despite these successes, applying self-play to enhance language model reasoning remains largely unexplored. Prior attempts have been limited to simple word games with offline updates (Cheng et al., 2024), LoRA adaptations that constrain learning capacity (Dettmers et al., 2023; Park et al., 2025), or single-turn code generation tasks (Zhao et al., 2025), falling short of leveraging multi-turn competitive dynamics that enable extended strategic reasoning.
@@
We introduce SPIRAL, which applies self-play to two-player zero-sum language games for developing reasoning capabilities. Zero-sum language games provide ideal training environments: they require strategic thinking and planning while remaining simple enough to enable stable learning, with clear rules that make outcomes verifiable. SPIRAL offers two key advantages. First, unlike traditional RLVR approaches that depend on human-curated problem-answer pairs, it generates infinite training data through game dynamics alone. Second, compared to fixed-opponent training (see Fig. 2), self-play prevents overfitting to static strategies by continuously evolving the challenge level. However, implementing this vision for LLMs presents significant technical challenges. The computational demands of multi-turn, multi-agent autoregressive generation require sophisticated distributed systems, while standard RL algorithms suffer from high variance in multi-agent settings. To address these challenges, we implement a fully online, multi-turn, multi-agent reinforcement learning system with a distributed actor-learner architecture. We also introduce role-conditioned advantage estimation (RAE), which stabilizes training by normalizing rewards relative to each player’s expected performance. Without RAE, our ablation experiments show that models suffer from thinking collapse, progressively abandoning reasoning traces that are critical for generalization.
@@
Key Findings. Training on zero-sum games produces reasoning capabilities that transfer broadly. SPIRAL trained solely on Kuhn Poker improves Qwen3-4B-Base by 8.6% on mathematical reasoning and 8.4% on general reasoning, outperforming SFT on 25,000 expert game trajectories despite using only game states with no mathematical content. This transfer works through three cognitive patterns we identify: systematic decomposition, expected value calculation, and case-by-case analysis. These patterns develop effectively through self-play, as fixed-opponent training fails while self-play continuously improves via adaptive challenge. Building on this, different games develop complementary skills (TicTacToe for spatial reasoning, Kuhn Poker for probabilistic thinking, and Simple Negotiation for strategic optimization) that combine synergistically in multi-game training. All these benefits require our Role-conditioned Advantage Estimation (RAE); without it, models abandon reasoning after 200 steps. The approach generalizes beyond weak models, with SPIRAL improving even strong models like DeepSeek-R1-Distill-Qwen-7B by 2.0%.
These results demonstrate that zero-sum games can serve as powerful reasoning gymnasiums, developing cognitive skills that transfer far beyond their training domain. Building on these findings, our work makes the following contributions:
@@
  1. 1.
    A Fully Online, Multi-Turn, Multi-Agent RL Framework for LLMs: We develop a distributed actor-learner architecture that enables online self-play with full-parameter updates across multiple two-player zero-sum language games. The multi-turn aspect trains models to reason through sequential decisions, directly preparing them for complex multi-step problem solving. Unlike prior offline approaches, this provides a continuous curriculum as the model must adapt to an ever-improving opponent: itself. We release our implementation111 https://github.com/spiral-rl/spiral. to facilitate further research.
  2. 2.
    Role-conditioned Advantage Estimation (RAE): We introduce a variance-reduced advantage estimator specifically designed for multi-agent settings. By normalizing rewards relative to each player’s expected performance, RAE reduces variance in multi-agent learning. Our ablation studies empirically validate that RAE stabilizes learning and prevents thinking collapse. Without it, models progressively abandon reasoning traces, which is critical for generalization.
  3. 3.
    Empirical Discovery of Transfer: We demonstrate that self-play on zero-sum games improves both out-of-distribution game performance and academic reasoning benchmarks without domain-specific training data. Our analysis identifies reasoning patterns (systematic decomposition, expected value calculation, case-by-case analysis) that transfer from games to mathematics at measurable rates. Different games develop specialized skills that transfer to related domains, and multi-game training combines these abilities synergistically.
@@ 2 Related Work
Reinforcement Learning for LLM Reasoning. Reinforcement learning has evolved from alignment tasks using RLHF (Jaques et al., 2019; Ouyang et al., 2022; Bai et al., 2022) to directly improving reasoning capabilities. Recent models like OpenAI o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek Team, 2024) demonstrate that RL with verifiable rewards (RLVR) can unlock chain-of-thought reasoning using rule-based rewards (Lightman et al., 2023; Uesato et al., 2022). However, these approaches depend on human-curated problem sets and domain-specific reward engineering. SPIRAL eliminates this dependency by using self-play games to generate infinite reasoning challenges without human supervision.
Multi-Agent RL for Language Models. Implementing MARL for full-scale LLMs presents significant technical challenges (Wan et al., 2025; Liu et al., 2025a). Prior work has made various compromises: Sarkar et al. (2025) uses RNNs instead of transformers; Jacob et al. (2022) focuses on simplified environments that don’t require full autoregressive generation; and Liao et al. (2024) shows the effectiveness of self-play with SFT on proprietary models. In contrast, SPIRAL implements fully online, full-parameter MARL through a distributed actor-learner architecture, enabling continuous adaptation to evolving opponents across multiple games.
Self-Play for Autonomous Improvement. Self-play in LLMs began with alignment methods like SPIN (Chen et al., 2024) and Self-Rewarding Language Models (Yuan et al., 2024). Recent work explores capability improvement: SPAG (Cheng et al., 2024) applies self-play to Adversarial Taboo but uses offline updates and remains confined to a single word game; SPC (Chen et al., 2025) and Genius (Xu et al., 2025) require human task distributions; Absolute Zero (Zhao et al., 2025) generates coding tasks but needs deterministic verification. SPIRAL demonstrates that self-play on strategic games improves academic benchmarks by 8.7% average without domain-specific data.
LLMs in Gaming. Games serve as testbeds for evaluating LLM capabilities (Paglieri et al., 2024; Ruoss et al., 2024; Zhang et al., 2024; Duan et al., 2024), but recent work explores games as training domains (Feng et al., 2024). LMRL-Gym (Abdulhai et al., 2023) benchmarks RL algorithms for LLMs in game environments; RAGEN (Wang et al., 2025) uses single-agent multi-turn RL; ViGaL (Xie et al., 2025) shows arcade games improve math reasoning; Divide-Fuse-Conquer (Zhang et al., 2025) trains on grouped games using offline learning. SPIRAL uniquely combines: (1) multi-agent self-play where both players share parameters, (2) fully online learning with continuous opponent evolution, and (3) transfer from zero-sum language games to reasoning without seeing any benchmark-related problems during training.
3 Preliminaries
Turn-level Markov Decision Process (MDP). Language model training traditionally formulates generation as a token-level MDP (Bellman, 1957; Rafailov et al., 2024) where each action is a single token from vocabulary $\mathcal{V}$. For multi-turn reasoning and game-playing, we instead adopt a turn-level MDP formulation $\mathcal{M}=(\mathcal{S},\mathcal{A},T,r,\gamma)$. Here, states $\mathcal{S}$ represent complete contexts (e.g., game configurations, problem states, or conversation histories), actions $\mathcal{A}$ are complete responses (containing many tokens), the transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ determines state dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ provides immediate rewards, and $\gamma\in[0,1]$ is the discount factor. The return is defined as the discounted sum of rewards: $R(\tau)=\sum_{t=0}^{T}\gamma^{t}r_{t}$.
where $c_{t}$ externalizes reasoning and $a_{t}\in\mathcal{A}$ is the executable action. (See App. A for how existing SFT and RLVR paradigms adapt to turn-level MDPs.)
Two-Player Zero-Sum Markov Games. We extend the single-agent MDP to competitive settings with a two-player zero-sum Markov game (Littman, 1994) $\mathcal{G}=(\mathcal{S},\mathcal{A}_{0},\mathcal{A}_{1},T,r,\gamma)$, where $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are the action spaces for player 0 and player 1 respectively. The zero-sum property requires:
@@ 4 The SPIRAL Framework
We present SPIRAL, a framework that enables language models to develop generalizable reasoning capabilities through multi-turn competitive self-play. The framework is illustrated in Fig. 3.
4.1 SPIRAL as Self-Play on Turn-Based Zero-Sum Games
SPIRAL implements self-play through turn-based zero-sum language games from a collection $\mathcal{G}=\{G_{1},G_{2},...,G_{n}\}$. Each game $G_{i}$ is a two-player zero-sum Markov game, with the specific structure that players alternate turns rather than acting simultaneously. SPIRAL leverages three properties of these games:
First, the multi-turn structure mirrors sequential reasoning problems. Players alternate turns: at time $t$, player $p=t\bmod 2$ acts while the opponent waits, formally: $a_{t}^{(p)}\in\mathcal{A}_{p}$ and $a_{t}^{(1-p)}=\emptyset$. Unlike single-turn bandit settings, this framework trains models to maintain context, plan ahead, and adapt strategies based on evolving game states.
Second, the zero-sum competitive dynamics create pressure for continuous improvement. Each game uses sparse terminal rewards: $r_{i}(s_{t},a_{t}^{(0)},a_{t}^{(1)})=0$ for all non-terminal states. At terminal state $s_{T}$, players receive opposite rewards:
$R_{0}(\tau)=\rho_{i}(s_{T}),\quad R_{1}(\tau)=-\rho_{i}(s_{T}),$ (3)
where $\rho_{i}:\mathcal{S}_{i}^{\text{terminal}}\rightarrow\{-1,0,1\}$ determines the outcome. This opposition forces models to develop robust strategies that cannot be easily exploited.
Finally, both players share a single policy $\pi_{\theta}$ with role conditioning. To enable this in the LLM setting, we prepend role-specific system prompts that specify whether the model is playing as Player 0 or Player 1 (see Sec. B.1 for example observations). At each turn, the active player generates:
$y_{t}^{(p)}\sim\pi_{\theta}(\cdot|s_{t},p,G_{i}),$ (4)
where conditioning on $p$ is achieved through the system prompt [^2] We note that both $p$ and $G_{i}$ can be subsumed into state $s_{t}$, but we choose to make them explicit here for clarity.. From $y_{t}^{(p)}$, we extract the action $a_{t}^{(p)}$ to update the game state. This shared-parameter approach ensures that as the model improves at one role, it simultaneously faces a stronger opponent, creating an automatic curriculum.
4.2 Multi-Turn Self-Play Training Objective
In self-play, we could train two separate policies $\pi_{\theta_{0}}$ and $\pi_{\theta_{1}}$ for each player. Each would maximize their own expected return:
$J_{0}(\theta_{0})$ $=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta_{0}}\times\pi_{\theta_{1}}|G}\left[R_{0}(\tau)\right]$ (5)
$J_{1}(\theta_{1})$ $=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta_{0}}\times\pi_{\theta_{1}}|G}\left[R_{1}(\tau)\right]$ (6)
However, SPIRAL uses a single shared policy $\pi_{\theta}$ for both players, setting $\theta_{0}=\theta_{1}=\theta$. While this creates opposing objectives since $R_{1}(\tau)=-R_{0}(\tau)$ in zero-sum games, the role conditioning through system prompts enables the model to learn distinct strategies for each position within a unified policy framework.
@@ 4.2 Multi-Turn Self-Play Training Objective
To optimize this shared policy, we apply policy gradient methods. Using REINFORCE (Williams, 1992), the gradient becomes:
@@
$\nabla_{\theta}J(\theta)=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta}\times\pi_{\theta}|G}\left[\underbrace{\sum_{t\in T_{0}}\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(0)}|s_{t},0,G)\cdot R_{0}(\tau)}_{\text{Player 0 gradient}}+\underbrace{\sum_{t\in T_{1}}\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(1)}|s_{t},1,G)\cdot R_{1}(\tau)}_{\text{Player 1 gradient}}\right],$ (7)
@@
where $T_{p}=\{t:t\bmod 2=p\}$ denotes turns where player $p$ acted. This formulation uses Monte Carlo returns which suffer from high variance, particularly problematic in self-play where the opponent’s strategy continuously evolves, making the environment highly non-stationary.
@@
9:              $p\leftarrow t\bmod 2$triangleright Determine active player
10:              $y_{t}^{(p)}\sim\pi_{\theta}(\cdot|s_{t},p,G_{i})$triangleright Generate reasoning + action
@@
12:              $a_{t}^{(1-p)}\leftarrow\emptyset$triangleright Inactive player
@@
16:         Add $(\tau,G_{i})$ to batch $\mathcal{B}$triangleright Store trajectory with its game
17:     end for
18:     // Role-Balanced Policy Optimization
19:     for $(\tau,G_{i})\in\mathcal{B}$ do
20:         for $p\in\{0,1\}$ do
21:              $b_{G_{i},p}\leftarrow\alpha b_{G_{i},p}+(1-\alpha)R_{p}(\tau)$
22:              $A_{G_{i},p}(\tau)\leftarrow R_{p}(\tau)-b_{G_{i},p}$
23:         end for
@@
25:     Update $\theta$ using policy gradient with advantages $A_{G_{i},p}$ (Eq. 10)
26:end while
4.3 Policy Optimization with Role-conditioned Advantage Estimation
To reduce the high variance inherent in multi-agent REINFORCE, we introduce Role-conditioned Advantage Estimation (RAE). In two-player games, even with a shared policy, different roles may have different expected returns due to game asymmetries (e.g., first-move advantage in TicTacToe, information asymmetry in Kuhn Poker).
RAE maintains separate baselines $b_{G,p}$ for each game $G\in\mathcal{G}$ and role $p\in\{0,1\}$, where each baseline estimates the expected return $\mathbb{E}[R_{p}(\tau)]$ for that role in that game. We update these baselines using exponential moving average (EMA) with a decay rate $\alpha\in[0,1]$:
$b_{G,p}\leftarrow\alpha b_{G,p}+(1-\alpha)R_{p}(\tau)$ (update baseline) (8)
$A_{G,p}(\tau)=R_{p}(\tau)-b_{G,p}$ (compute advantage) (9)
This provides better variance reduction than a global baseline by accounting for role-specific asymmetries, such as the first-move advantage. The variance-reduced policy gradient becomes:
@@ 4.3 Policy Optimization with Role-conditioned Advantage Estimation
By centering returns around role-specific expectations, RAE ensures that gradient updates reflect genuine learning signal rather than inherent advantages of certain player positions. Note that we do not normalize by response length to avoid length bias (Liu et al., 2025b). The complete procedure is in Algorithm 1.
4.4 Implementation

### v3

# arXiv:2506.24119v3[cs.AI] 02 Mar 2026

@@ Abstract
Recent advances in reinforcement learning have shown that language models can develop sophisticated reasoning through training on tasks with verifiable rewards, but these approaches depend on human-curated problem-answer pairs and domain-specific reward engineering. We introduce SPIRAL, a self-play framework where models learn by playing multi-turn, zero-sum games against continuously improving versions of themselves, generating an automatic curriculum of stronger opponents, and eliminating the need for human supervision. To enable this self-play training at scale, we implement a fully online, multi-turn, multi-agent reinforcement learning system for LLMs and propose role-conditioned advantage estimation (RAE) to stabilize multi-agent training. SPIRAL produces reasoning capabilities that transfer broadly, improving performance by up to 10% across a suite of 8 reasoning benchmarks on 4 different models spanning Qwen and Llama model families, outperforming supervised fine-tuning on 25,000 expert game trajectories. Multi-game training (TicTacToe, Kuhn Poker, Simple Negotiation) yields the strongest results, with improvements observed across both base and instruction-tuned models. Analysis of chain-of-thought traces reveals that games develop distinct cognitive patterns that transfer to improve reasoning performance, with different games developing complementary strengths. Even models which have already been trained on reasoning tasks using RLVR, like DeepSeek-R1-Distill-Qwen-7B, still benefit from our approach. These results demonstrate that zero-sum games naturally develop transferable reasoning capabilities across diverse model architectures and training stages, highlighting a promising direction for autonomous reasoning development. Our code can be found in https://github.com/spiral-rl/spiral.
@@ 1 Introduction
However, current approaches face a fundamental scalability bottleneck: dependence on carefully engineered reward functions, domain-specific datasets, and expert supervision (DeepSeek Team, 2024; Ouyang et al., 2022; Bai et al., 2022). Each new reasoning domain requires experts to craft evaluation metrics, curate training problems, and validate reasoning traces. This manual process becomes increasingly unsustainable as we pursue more general intelligence, limiting both scale and diversity of reasoning challenges that models can learn from.
@@
Self-play on games offers a solution by eliminating human supervision in training data creation (Silver et al., 2017; Tesauro, 1995). In game-based self-play, models learn by competing against copies of themselves, where game outcomes provide automatic feedback and opponents improve equally, maintaining a consistent challenge that drives continuous learning. Although many prominent successes in past AI research relied on self-play—from TD-Gammon (Tesauro, 1995) to AlphaGo (Silver et al., 2016; 2017) to OpenAI Five (Berner et al., 2019)—so far, applying self-play on games to enhance language model reasoning remains largely unexplored. Prior attempts have been limited to simple word games with offline updates (Cheng et al., 2024), LoRA adaptations (Dettmers et al., 2023; Park et al., 2025), or single-turn tasks (Zhao et al., 2025), falling short of leveraging multi-turn competitive dynamics for extended strategic reasoning.
@@
We introduce SPIRAL (Self-Play on zero-sum games Incentivizes Reasoning via multi-Agent multi-turn reinforcement Learning), which applies self-play to two-player zero-sum language games for developing reasoning capabilities. SPIRAL offers two key advantages: unlike traditional RLVR approaches depending on human-curated problem-answer pairs, it generates unlimited training data through game dynamics alone; compared to fixed-opponent training (see Fig.˜2), self-play prevents overfitting to static strategies by continuously evolving challenge level. However, implementing this for LLMs presents significant challenges. The computational demands of multi-turn, multi-agent autoregressive generation require sophisticated distributed systems, while standard RL algorithms suffer from high variance in multi-agent settings. We address these through a fully online, multi-turn, multi-agent reinforcement learning system with distributed actor-learner architecture and introduce role-conditioned advantage estimation (RAE), which stabilizes training by normalizing rewards relative to each player’s expected performance.
@@
Key Findings. Training on zero-sum games produces reasoning capabilities that transfer broadly across diverse model architectures. Multi-game SPIRAL training (TicTacToe, Kuhn Poker, Simple Negotiation) achieves up to 10% improvement across 8 reasoning benchmarks, outperforming supervised fine-tuning on 25,000 expert trajectories. On Qwen3-4B-Base (Yang et al., 2025), multi-game training reaches 44.5% average performance versus 34.0% baseline (+10.5+10.5% absolute gain), while Qwen3-8B-Base (Yang et al., 2025) improves from 39.5% to 49.6% (+10.1+10.1%). The approach generalizes across model families: base models (Qwen3-4B/8B-Base, Octothinker-8B-Base; Wang et al. (2025a)) and instruction-tuned models (Llama-3.1-8B-Instruct; Dubey et al. (2024)) all show consistent improvements, with Octothinker-8B-Base gaining 8.0% and Llama-3.1-8B-Instruct improving 2.0% despite already being instruction-tuned. Each game develops complementary cognitive skills: TicTacToe for spatial reasoning, Kuhn Poker for probabilistic thinking, and Simple Negotiation for strategic optimization, which combine synergistically in multi-game training. Using post-hoc analysis, we find examples of three patterns learned from gameplay that transfer to improve math performance: case-by-case analysis, expected value calculation, and pattern recognition. These patterns develop effectively through self-play’s adaptive curriculum, as fixed-opponent training fails while self-play continuously improves. Role-conditioned Advantage Estimation proves critical: without RAE, models abandon reasoning after 200 steps, progressively generating empty thinking traces that destroy generalization. Building on these findings, our work makes the following contributions:
  1. 1.
    A Fully Online, Multi-Turn, Multi-Agent RL Framework for LLMs: We develop a distributed actor-learner architecture enabling online self-play with full-parameter updates across multiple two-player zero-sum language games. The multi-turn aspect trains models to reason through sequential decisions, directly preparing them for complex multi-step problem solving. Unlike prior offline approaches, this provides continuous curriculum as the model adapts to an ever-improving opponent. We release our implementation to facilitate further research.
  2. 2.
    Role-conditioned Advantage Estimation (RAE): We introduce a variance-reduced advantage estimator specifically designed for multi-agent settings. By normalizing rewards relative to each player’s expected performance, RAE prevents the degradation of the model’s reasoning capabilities, a failure mode we term “thinking collapse”. Without it, models progressively abandon reasoning traces after 200 steps, which is critical for generalization.
  3. 3.
    Empirical Discovery of Transfer: We demonstrate that self-play on zero-sum games improves both out-of-distribution game performance and academic reasoning benchmarks by up to 10% without domain-specific training data. Our analysis identifies reasoning patterns (systematic decomposition, expected value calculation, case-by-case analysis) that transfer from games to mathematics at measurable rates, with different games developing specialized skills that combine synergistically in multi-game training.
@@ 2 Related Work
Reinforcement Learning for LLM Reasoning. Reinforcement learning (RL) in LLMs has progressed from alignment tasks using RLHF (Jaques et al., 2019; Ouyang et al., 2022; Bai et al., 2022) to directly improving reasoning capabilities. Recent models like OpenAI o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek Team, 2024) demonstrate that RL with verifiable rewards (RLVR) can unlock chain-of-thought reasoning using rule-based rewards (Lightman et al., 2023; Uesato et al., 2022). However, these approaches depend on human-curated problem sets and domain-specific reward engineering. SPIRAL eliminates this dependency by using self-play games to generate unlimited reasoning challenges without human supervision.
Self-Play and Multi-Agent RL for LLMs. Self-play in LLMs initially focused on alignment objectives (Chen et al., 2024; Yuan et al., 2024) before recent work applied it to enhance model capabilities. SPAG (Cheng et al., 2024) applies self-play to Adversarial Taboo using offline updates on a single game; SPC (Chen et al., 2025) and Genius (Xu et al., 2025a) require predefined human task distributions; Absolute Zero (Zhao et al., 2025) generates single-turn coding tasks; Foundation Model Self-Play (Dharna et al., 2025) uses foundation models to evolve code-based policies rather than direct gameplay; Prover-Verifier Game (Kirchner et al., 2024) improve output legibility through adversarial training. Implementing multi-agent RL (MARL) for full-scale LLMs presents significant technical challenges (Wan et al., 2025; Liu et al., 2025b; a). Prior work circumvents these challenges by using RNNs instead of transformers (Sarkar et al., 2025), restricting to simplified environments (Jacob et al., 2022; Sukhbaatar et al., 2018), or applying supervised fine-tuning on trajectories from proprietary models (Liao et al., 2024). In contrast, SPIRAL implements fully online, full-parameter MARL through a distributed actor-learner architecture, enabling continuous adaptation to evolving opponents across multiple games.
LLMs in Gaming. Games serve as both evaluation benchmarks (Paglieri et al., 2024; Ruoss et al., 2024; Zhang et al., 2024; Duan et al., 2024) and training domains (Feng et al., 2024; Verma et al., 2025). LMRL-Gym (Abdulhai et al., 2023) and RAGEN (Wang et al., 2025b) both employ single-agent multi-turn RL, with LMRL-Gym providing 8 benchmarking tasks and RAGEN focusing on trajectory-level optimization. ViGaL (Xie et al., 2025b) shows that single-agent RL on visual-spatial games transfers to mathematical reasoning without explicit math training. Logic-RL (Xie et al., 2025a) trains on synthetic puzzle games; Divide-Fuse-Conquer (Zhang et al., 2025) applies offline learning to grouped games; Boundless Socratic Learning (Schaul, 2024) uses language games for continual learning; Code2Logic (Tong et al., 2025) synthesizes reasoning data from game code. We also distinguish SPIRAL from works designed for achieving super-human performance at specific games. Cicero (FAIR et al., 2022) integrates a language model with a separate strategic planning algorithm to achieve human-level performance in Diplomacy. Similarly, agents developed for Werewolf (Xu et al., 2024; 2025b) and Avalon (Wang et al., 2023) focus on optimizing strategic communication and hidden-role deduction to win within their respective game rules. While these works target in-domain victory, SPIRAL treats competitive pressure as a training scaffold to develop reasoning patterns (e.g., case analysis) that transfer to out-of-domain reasoning tasks. SPIRAL uniquely combines three elements: (1) multi-agent self-play where both players share parameters, (2) fully online learning with continuous opponent evolution, and (3) demonstrated transfer from zero-sum language games to academic reasoning benchmarks achieving up to 10.5% improvement without exposure to benchmark-related problems during training.
@@ 3 The SPIRAL Framework
We present SPIRAL, a framework enabling language models to develop generalizable reasoning through multi-turn competitive self-play on games, illustrated in Figure 3.
Formulation. SPIRAL implements self-play through turn-based zero-sum language games from collection $\mathcal{G}=\{G_{1},G_{2},...,G_{n}\}$. Each game $G_{i}$ is a two-player zero-sum Markov game (Littman, 1994) built on turn-level MDPs where states $s\in\mathcal{S}$ represent complete contexts (e.g., game configurations), actions $a\in\mathcal{A}$ are complete multi-token responses, and transition function $T_{i}$ determines state dynamics after full turn completion. The zero-sum property ensures $r_{0}(s,a^{(0)},a^{(1)})+r_{1}(s,a^{(0)},a^{(1)})=0$ for all states and actions where $a^{(p)}$ denotes the action of player $p\in\{0,1\}$, creating competitive dynamics. See Appendix C for detailed formulations.
Benefits of multi-turn, zero-sum games. Zero-sum dynamics create continuous improvement pressure through rewards given only at game termination: $r_{i}(s_{t},a_{t}^{(0)},a_{t}^{(1)})=0$ for all non-terminal states, with terminal rewards $R_{0}(\tau)=\rho_{i}(s_{T})$ and $R_{1}(\tau)=-\rho_{i}(s_{T})$ where $\rho_{i}:\mathcal{S}_{i}^{\text{terminal}}\rightarrow\{-1,0,1\}$ determines the outcome and $\tau$ represents the complete trajectory. This forces robust strategy development as models only receive feedback upon game completion. The multi-turn structure mirrors sequential reasoning problems: players alternate turns with $p=t\bmod 2$ acting at time $t$ while the opponent waits, training models to maintain context, plan ahead, and adapt strategies.
Self-play. Rather than training separate policies $\pi_{\theta_{0}}$ and $\pi_{\theta_{1}}$ for each player, SPIRAL uses a single shared policy $\pi_{\theta}$ with parameters $\theta$, setting $\theta_{0}=\theta_{1}=\theta$. Role conditioning through system prompts enables the model to learn distinct strategies for each position (see Appendix D.1). At each turn, the active player generates a complete response $y_{t}^{(p)}\sim\pi_{\theta}(\cdot|s_{t},p,G_{i})$ conditioned on current state $s_{t}$, player role $p$, and game $G_{i}$. From this response, we extract the action $a_{t}^{(p)}$ to update the game state via $s_{t+1}=T_{i}(s_{t},a_{t}^{(0)},a_{t}^{(1)})$ where $a_{t}^{(1-p)}=\emptyset$ for the inactive player. This shared-parameter approach ensures efficient use of GPU memory while also guaranteeing that as the model improves at one role, it simultaneously faces a stronger opponent, creating an automatic curriculum. Algorithm 1 presents the complete training procedure.
@@
RL objective. To optimize this shared policy, we apply Monte Carlo policy gradient methods. Using REINFORCE (Williams, 1992), the gradient becomes:
@@
$\nabla_{\theta}J(\theta)=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta}\times\pi_{\theta}|G}\left[\sum_{t\in T_{0}}\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(0)}|s_{t},0,G)\cdot R_{0}(\tau)+\sum_{t\in T_{1}}\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(1)}|s_{t},1,G)\cdot R_{1}(\tau)\right],$ (1)
@@
where $T_{p}=\{t:t\bmod 2=p\}$ denotes turns where player $p$ acted. This formulation uses Monte Carlo returns which suffer from high variance, particularly problematic in self-play where the opponent’s strategy continuously evolves, making the environment non-stationary.
@@
9:     $p\leftarrow t\bmod 2$ $\triangleright$ Determine active player
10:     $y_{t}^{(p)}\sim\pi_{\theta}(\cdot|s_{t},p,G_{i})$ $\triangleright$ Generate ‘[reasoning] [action]
@@
12:     $a_{t}^{(1-p)}\leftarrow\emptyset$ $\triangleright$ Inactive player
@@
16:   Define $\tau=\{(s_{0},y_{0}),(s_{1},y_{1}),\dots,(s_{T},y_{T})\}$
17:   Add $(\tau,G_{i})$ to batch $\mathcal{B}$ $\triangleright$ Store trajectory with its game
18:  end for
19:  // Role-Balanced Policy Optimization
20:  for $(\tau,G_{i})\in\mathcal{B}$ do
21:   for $p\in\{0,1\}$ do
22:     $b_{G_{i},p}\leftarrow\alpha b_{G_{i},p}+(1-\alpha)R_{p}(\tau)$
23:     $A_{G_{i},p}(\tau)\leftarrow R_{p}(\tau)-b_{G_{i},p}$
@@
25:  end for
26:  Update $\theta$ on full sequences $y_{t}$ using REINFORCE with advantages $A_{G_{i},p}$ (Eq. 3)
27:end while
Role-conditioned advantage estimation. Self-play on zero-sum games implies using the same model to optimize for opposing objectives, since $R_{1}(\tau)=-R_{0}(\tau)$. This can lead to unstable training dynamics which impedes learning. To reduce the high variance inherent in multi-agent REINFORCE, we introduce Role-conditioned Advantage Estimation (RAE). In two-player games, even with a shared policy, different roles may have different expected returns due to game asymmetries (e.g., first-move advantage in TicTacToe, information asymmetry in Kuhn Poker). RAE maintains separate baselines $b_{G,p}$ for each game $G\in\mathcal{G}$ and role $p\in\{0,1\}$, estimating expected return $\mathbb{E}[R_{p}(\tau)]$ for that role in that game. We update these baselines using exponential moving average with decay rate $\alpha\in[0,1]$:
$b_{G,p}$ $\leftarrow\alpha b_{G,p}+(1-\alpha)R_{p}(\tau),\quad A_{G,p}(\tau)=R_{p}(\tau)-b_{G,p}$ (2)
This provides better variance reduction than a global baseline by accounting for role-specific asymmetries. The variance-reduced policy gradient becomes:
@@
By centering returns around role-specific expectations, RAE ensures gradient updates reflect genuine learning signal rather than inherent positional advantages. We do not normalize by response length to avoid length bias (Liu et al., 2025c). The complete procedure is in Algorithm 1.